New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is partially correct. On March 1, 2017, Sandhill Company sold 25,100 of its 9%, 20-year, $1,000 face value bonds at 96. Interest payment dates are March 1 and September 1, and the company uses the straight-line method of bond discount amortization. On February 1, 2018, Sandhill took advantage of favorable prices of its stock to extinguish 2,850 of the bonds by issuing 149,800 shares of its $1 par value common stock. At this time, the accrued interest was paid in cash. The company’s stock was selling for $20.75 per share on February 1, 2018.

Prepare the journal entries needed on the books of Sandhill Company to record the following. (a) March 1, 2017: issuance of the bonds. (b) September 1, 2017: payment of semiannual interest. (c) December 31, 2017: accrual of interest expense. (d) February 1, 2018: extinguishment of 2,850 bonds. (No reversing entries made.)
Answer #1
Date Account Titles Debit Credit
Mar 1, 2017 Cash $    24,096,000 =25100*960
Discount on Bonds Payable $      1,004,000 =25100*(1000-960)
      Bonds Payable $          25,100,000 =25100*1000
(Bonds Issued at discount)
Sep 1, 2017 Interest Expense $      1,154,600 =1129500+25100
         Cash $            1,129,500 =25100000*9%*6/12
        Discount on Bonds Payable $                  25,100 =1004000/40
(Interest paid and discount amortized)
Dec 31, 2017 Interest Expense $          769,733 =1154600*4/6
        Interest Payable $                753,000 =25100000*9%*4/12
        Discount on Bonds Payable $                  16,733 =1004000/40*4/6
(Interest expense and discount amortized)
Feb 1, 2017 Interest Expense $            21,850 =1154600/25100*2850*1/6
Interest Payable $            85,500 =753000*2850/25100
        Cash $                106,875 =2850*1000*9%*5/12
        Discount on Bonds Payable $                        475 =107350-106875
(Payment of Interest accrued on bonds redeemed)
Feb 1, 2017 Bonds Payable $      2,850,000 =2850*1000
Loss on Redemption of Bonds $          367,125 =108775+149800+2958550-2850000
        Discount on Bonds Payable $                108,775 =2850*40-4750-475
        Common Stock ($1 par value) $                149,800 149800
        Paid in Capital in excess of Par - Common Stock $            2,958,550 =149800*19.75

Discount already amortized on bonds extinguished = ($25100+16733)/25100*280 + 475 = $5225